# MotiF: Making Text Count in Image Animation with Motion Focal Loss

### v1

@@ Abstract
Text-Image-to-Video (TI2V) generation aims to generate a video from an image following a text description, which is also referred to as text-guided image animation. Most existing methods struggle to generate videos that align well with the text prompts, particularly when motion is specified. To overcome this limitation, we introduce MotiF, a simple yet effective approach that directs the model’s learning to the regions with more motion, thereby improving the text alignment and motion generation. We use optical flow to generate a motion heatmap and weight the loss according to the intensity of the motion. This modified objective leads to noticeable improvements and complements existing methods that utilize motion priors as model inputs. Additionally, due to the lack of a diverse benchmark for evaluating TI2V generation, we propose TI2V Bench, a dataset consists of $320$ image-text pairs for robust evaluation. We present a human evaluation protocol that asks the annotators to select an overall preference between two videos followed by their justifications. Through a comprehensive evaluation, MotiF outperforms nine open-sourced models, achieving an average preference of $72\%$. The TI2V Bench is released in the project website.
Figure 1: Motivation and results of MotiF. (a) Example video frames and the corresponding motion heatmaps calculated from optical flow. In this example, $97\%$ of the pixels are static while only $3\%$ has meaningful motion. (b) In standard TI2V training pipeline, the model may learn to over-rely on the conditional image to optimize the L2 loss. This issue has been identified in [53] and termed as conditional image leakage. We propose MotiF to guide the model’s learning to focus on regions with more motion via motion heatmap re-weighting. (c) Qualitative results comparing MotiF to the baseline on examples from our proposed TI2V Bench.
@@ 1 Introduction
From creating short clips of moving MNIST digits [27] to producing a minute-long high-definition video featuring a stylish walking woman [5], Text-to-Video (T2V) generation [18, 37, 3, 30] has progressed remarkably over the past decade. Meanwhile, video generation with additional conditions [22, 31, 15, 43, 45, 44] has also gained considerable attention. In this paper, we focus on Text-Image-to-Video (TI2V) generation, a task first introduced in [19] aiming to generate a video based on an initial image and a text description [46, 50, 53, 26]. Our goal is to enhance text adherence in open-domain TI2V generation. Commonly known as Image-to-Video (I2V) generation  [50, 53] or image animation [46, 25], this task often overlooks the text guidance. In many cases, text merely describes the image or video [21, 13], or is unnecessary [49, 23, 6], underscoring the need for focused improvement.
@@
The main challenge in TI2V modeling is that while the image offers a strong spatial signal to learn the video content, the model must rely mainly on the text to infer motion. When given an image and multiple different prompts, most prior works generate videos with limited motion. Zhao et al. [53] identified this issue, termed as conditional image leakage, where the model overly relies on the image condition. To address this issue, some studies propose to reduce the strength of the image condition by adding noise or masking the image [53, 26]. Others seek to derive more motion priors to facilitate motion learning [10, 51]. These methods all emphasis on enhancing the input signals, hoping that the model can learn to leverage them implicitly. In contrast, we propose to improve the training objective to enable the model to focus on the regions with more motion explicitly, which is an orthogonal direction to prior works and can be easily combined with existing techniques.
@@
We hypothesize that the model’s difficulty in following motion-based instructions may stem from insufficient attention to motion patterns. As shown in Figure 1 (a), in a video with a static background, $97\%$ of the pixels remain unchanged over time, while only $3\%$ exhibit meaningful motion. This subtle motion pattern might be neglected by the model especially considering that all regions are treated equally in the L2 loss. Motivated by this, we introduce Motion Focal loss (MotiF), a term inspired from the focal loss in dense object detection [35]. Specifically, we first use optical flow [42] to create a motion heatmap that represents the motion intensities of the video. We then use the motion heatmap to assign loss weight for the video to focus on regions with more motion. Additionally, we analyze commonly used image conditioning mechanisms and observe that concatenating the conditioning image with the diffusion input works best for both image and text alignment.
@@
Beyond training, evaluating TI2V generation is also challenging due to the lack of a suitable evaluation set and reliable metrics. Existing evaluation sets are limited, they are either too small [51] or the text is used as a description of the starting image rather than specifying the intended motion for animation [13]. Therefore, we introduce a new benchmark, TI2V Bench, that consists of $320$ image-text pairs covering $22$ different scenarios. Each scenario includes $3$ to $5$ images with similar contents in different styles, and $3$ to $5$ prompts to animate these images to produce different motion. We intentionally include challenging scenarios such as novel object introduction and fine-grained object references. We use the Emu model [11], publicly available through meta.ai, to generate the initial images.
@@
Previous studies have employed various automatic metrics to assess different quality aspects, such as video quality, motion intensity, and text alignment. We observed that these metrics may not always align with human perception. Following Movie Gen [30], we mainly rely on human evaluation with A-B testing to evaluate the performance. Our human evaluation protocol is motivated by the JUICE protocol [15], where annotators are asked to indicate an overall preference and justify their choices across several axes, including image alignment, text alignment, object motion, and overall quality. Through our comprehensive human evaluations, we show that MotiF enhances TI2V generation capabilities, notably improving text alignment and motion quality. In summary, we make the following contributions:
@@
  • We present Motion Focal Loss (MotiF) that encourages TI2V generation to concentrate on regions with larger motion, and is complementary to existing techniques.
@@
  • We present a new benchmark, TI2V Bench, including synthetic images and texts that cover a wide range of complex scenarios, as well as a human evaluation protocol designed to assess TI2V generation performance.
@@ 2.1 Text-Image-to-Video Generation
With the rapid advancements in Text-to-Image (T2I) and Text-to-Video (T2V) generation, TI2V generation has also gained significant attention. Although training-free methods [49, 28] exist that utilize pretrained T2I or T2V models with inference optimization to enable zero-shot TI2V generation, we focus on approaches that require training, especially on the most widely adapted latent diffusion models [1, 54, 37]. The techniques in prior TI2V generation methods can be mainly grouped into two categories: 1) design novel architectures to integrate the image condition, 2) derive motion priors to improve motion learning. Our approach falls into the second category.
Image Condition Integration. Most recent works use either the U-Net architecture [18, 17] or the Diffusion Transformer (DiT) [29] for video generation, where the image condition can be incorporated similarly. Most works integrate the image by concatenating it with the noisy video [15, 3, 25, 10], which serves as a spatial-aligned content guidance. Additionally, DynamiCrafter [46], ConsistI2V [33] and LivePhoto [10] also use cross-attention layers on the image embeddings to further strengthen the image condition and thus improve consistency in video generation. On the contrary, Zhao et al. [53] applied higher noise levels at later time steps to weaken the image condition to tackle the condition image leakage issue. Similarly, Follow-Your-Click [26] introduced a first-frame masking strategy to degrade the image condition. In our work, instead of disrupting the image condition to shift the model’s focus on the text, our training objective explicitly encourages the model to focus on video regions with more motion, and thus be more faithful to the text description.
Motion Priors. Our world exhibits a rich diversity of unique motion patterns. Modeling motion from real-world videos for generation is challenging due to the inherent ambiguity that arises when a single description can correspond to multiple possible motions. Many methods explore motion priors to improve the motion learning. A common approach is to calculate a motion score [10, 12, 26, 51, 31, 25, 16] from the video and embed it into the model, helping it to more easily learn the motion level. It also allows for controlling motion intensity during inference. Previous works primarily differ in how they calculate the motion score, yet they generally follow a similar concept based on frame similarities. Additionally, LivePhoto [10] introduced a framework to re-weight text embeddings to place greater emphasis on the motion-related words. Cinemo [25] introduced a novel strategy to learn motion residuals, calculated by subtracting the first frame from each subsequent frame. MoCA[48] introduces optical flow as an additional conditioning for image animation. Similar to our work, Follow-Your-Click [26] also used optical flow to calculate a mask, but the mask is embedded into the model and thus it requires additional input from the user during inference time. In our case, the motion prior is leveraged at the training objective level, which is explicit, simple and does not require any additional input during inference.
@@ 2.2 Video Generation Evaluation
Benchmarks and Automatic Metrics. Make-a-Video is the first to report FVD/IS on the UCF-101 dataset [41] and FID/CLIP similarities on the MSR-VTT dataset [47] under the zero-shot setting, which is followed by later works in both T2V [4, 54, 1, 14] and TI2V generation [46, 33, 25, 53]. Re-purposing existing video understanding datasets for TI2V evaluation is undesirable due to their unconstrained data collection procedures, which may not capture the nuances of TI2V tasks. As a result, recent studies have begun to collect new evaluation sets [33, 13, 51] and explore additional automatic metrics for more comprehensive evaluation [24, 21, 52]. AIGCBench [13] is proposed for TI2V task. However, their text descriptions are used to generate the image but not to drive the motion generation, which neglected the core idea of text-guided image animation. AnimateBench [51] shares some similarities with our benchmark approach of using T2I models to generate images and applying different text prompts to animate the same image. However, the number of unique images and texts in AnimateBench are relatively small and the domains are also limited. Furthermore, the automatic metrics employed in their work have several drawbacks as shown in Table 4.
Human Evaluation. Following Movie Gen [30], we rely on human annotators to evaluate the performance of TI2V generation methods. Prior works used different human evaluation protocols including A-B testing [53], A-testing [10] and ranking multiple methods [26, 16, 51]. While A-testing and ranking are more efficient, we believe A-B testing is less ambiguous and may yield more reliable results. Unlike previous works that evaluate multiple metrics, we ask the annotators to select the best video and justify their choices based on various aspects, following the JUICE design [15]. This will generate a single metric for calibration while also enabling a detailed examination of results through the various quality justifications. To the best of our knowledge, we have conducted the most comprehensive human evaluation by comparing to nine open-sourced TI2V models and shown the advantage of MotiF.
@@ 3.2 Modeling
Figure 3: Example image-text pairs in TI2V Bench. For each scenario (column), we first think of a scene that could be potentially animated to generate different types of motion. We include challenging scenarios when there are multiple objects (yellow/blue/red balloon) in the initial image for fine-grained control or the text prompt describes a new object (frisbee, bubbles) to enter the scene. Then we come up with different prompts and use the publicly available meta.ai tool to generate diverse sets of images. Images of low quality or those not in the appropriate initial state are removed.
@@
TI2V Bench (Ours) image(synthetic)-text Yes $320$ $88$ $133$ human TI2V score
@@ 4.1 Evaluation Set
While it’s relatively easy to collect a diverse set of prompts for T2V evaluation, evaluating TI2V is not straightforward as it requires a conditional image. As shown in Table 1, recent TI2V benchmarks can be classified into three categories: 1) video-text pairs; 2) image-text pairs with realistic images; 3) image-text pairs with synthetic images. Re-purposing existing video understanding datasets [41, 47, 2] for TI2V evaluation is non ideal, as these datasets assume only one possible outcome from the initial frame, and the text descriptions typically do not emphasize the intended motion. Existing images are also limited in their diversity and potential motion changes. Thus, we resort to image-text pairs with synthetic images to curate TI2V Bench. AIGCBench [13] and Animate Bench [51] are most relevant to our work. However, the texts in AIGCBench are used to generate the images but not meant to describe the motion in the videos, which is critical for TI2V task. Animate Bench employs motion-based text descriptions, but its evaluation set is relatively small and limited, where the images often feature a single object or scene, and the text description lacks diversity.
Our goal is to curate an evaluation set that has a diverse set of images, and each image has a diverse set of prompts for text-guided animation. Some examples are shown in Figure 3. To create this benchmark, we first design scenarios in which multiple actions can be performed from an initial state and enumerate a range of possible actions. For example, in the scenario of “an aerial view of a car on the road,” possible actions could include “moving forward,” “moving backward,” and “changing its lane.” We come up with $22$ different scenarios with diverse objects and scenes, each with multiple options for different motion. We make sure to include challenging scenarios like images with multiple objects, and text descriptions that introduce a new object expected to appear in the video.
@@
Figure 4: Human evaluation results comparing MotiF to nine open-sourced models [46, 7, 25, 50, 9, 53, 28, 12, 33] on our proposed TI2V Bench. We achieved considerable improvements across the board with an average preference of $72\%$. Through examining the justification choices, we found that our model mostly excel at improving text alignment and object motion, which matches very well with our motivation. Note that all the inference of prior works are done at Brown University.
@@ 5 Experiment
Figure 5: Qualitative comparison to prior works on TI2V Bench. Sampled frames are ordered from left to right.
@@ 5.2 Comparison to Prior Works
We compare to nine open-sourced TI2V generation methods. To the best of our knowledge, this is the most comprehensive TI2V human evaluation. We ensure fair comparisons by following each method’s pre-processing and post-processing pipeline and ensure the generated videos are natural after resizing to the same resolution $(320\times 320)$. The results are shown in Figure 4. Our method wins by a considerable margin to prior works. From the justification selections, the main reasons for our method to win are on the object motion and text alignment, which is exactly the motivation of MotiF to improve text-driven motion learning.
@@
In addition to human evaluation, we also report the automatic evaluation on Animate Bench [51] that is most relevant to TI2V Bench. As shown in Table 4, MotiF achieved comparable performance to prior works. It’s worth noting that a static video baseline (by repeating the condition image) gets the best Image Alignment score and reasonable Text Alignment score, suggesting that these metrics are non ideal for evaluating TI2V generation. This has motivated us to mainly rely on human evaluation.

### v2

@@ Abstract
Text-Image-to-Video (TI2V) generation aims to generate a video from an image following a text description, which is also referred to as text-guided image animation. Most existing methods struggle to generate videos that align well with the text prompts, particularly when motion is specified. To overcome this limitation, we introduce MotiF, a simple yet effective approach that directs the model’s learning to the regions with more motion, thereby improving the text alignment and motion generation. We use optical flow to generate a motion heatmap and weight the loss according to the intensity of the motion. This modified objective leads to noticeable improvements and complements existing methods that utilize motion priors as model inputs. Additionally, due to the lack of a diverse benchmark for evaluating TI2V generation, we propose TI2V-Bench, a dataset consists of $320$ image-text pairs for robust evaluation. We present a human evaluation protocol that asks the annotators to select an overall preference between two videos followed by their justifications. Through a comprehensive evaluation, MotiF outperforms nine open-sourced models, achieving an average preference of $72\%$. The TI2V-Bench and additional results are released in the project website.
Figure 1: Motivation and results of MotiF. (a) Example video frames and the corresponding motion heatmaps calculated from optical flow. In this example, $97\%$ of the pixels are static while only $3\%$ has meaningful motion. (b) In standard TI2V training pipeline, the model may learn to over-rely on the conditional image to optimize the L2 loss. This issue has been identified in [53] and termed as conditional image leakage. We propose MotiF to guide the model’s learning to focus on regions with more motion via motion heatmap re-weighting. (c) Qualitative results comparing MotiF to the baseline on examples from our proposed TI2V-Bench evaluation set.
@@ 1 Introduction
From creating short clips of moving MNIST digits [27] to producing a minute-long high-definition video featuring a stylish walking woman [5], Text-to-Video (T2V) generation [18, 37, 3, 30] has progressed remarkably over the past decade. Meanwhile, video generation with additional conditions [22, 31, 15, 43, 45, 44] has also gained considerable attention. In this paper, we focus on Text-Image-to-Video (TI2V) generation, a task first introduced in [19] aiming to generate a video based on an initial image and a text description [46, 50, 53, 26]. Our goal is to improve text adherence in open-domain TI2V generation. Commonly known as Image-to-Video (I2V) generation  [50, 53] or image animation [46, 25], this task often overlooks the text guidance. In many cases, text merely describes the image or video [21, 13], or is unnecessary [49, 23, 6], underscoring the need for focused improvement.
@@
The main challenge in TI2V modeling is that, while the image offers a strong spatial signal to learn the video content, the model must rely mainly on the text to infer motion. When given an image and multiple different prompts, most previous work generates videos with limited motion. Zhao et al. [53] identified this issue as conditional image leakage, where the model overly relies on the image condition. To address this issue, some studies propose to reduce the strength of the image condition by adding noise or masking the image [53, 26]. Others seek to derive more motion priors to facilitate motion learning [10, 51]. All of these methods are based on enhancing the input signals, in the hope that the model can learn to leverage them implicitly. In contrast, we propose to improve the training objective to enable the model to focus on the regions with more motion explicitly, which is an orthogonal direction to prior works and can be easily combined with existing techniques.
@@
We hypothesize that the model’s difficulty in following motion-based instructions may stem from insufficient attention to motion patterns. As shown in Figure 1 (a), in a video with a static background, $97\%$ of the pixels remain unchanged over time, while only $3\%$ exhibit meaningful motion. This subtle motion pattern might be neglected by the model especially considering that all regions are treated equally in the L2 loss. Motivated by this, we introduce Motion Focal loss (MotiF), a term inspired by the focal loss in dense object detection [35]. Specifically, we first use optical flow [42] to create a motion heatmap that represents the motion intensities of the video. We then use the motion heatmap to assign loss weight for the video to focus on regions with more motion. Additionally, we analyze commonly used image conditioning mechanisms and observe that concatenating the conditioning image with the diffusion input works best for both image and text alignment.
@@
Beyond training, evaluating TI2V generation is also challenging due to the lack of a suitable evaluation set and reliable metrics. Existing evaluation sets are limited, they are either too small [51] or the text is used as a description of the starting image rather than specifying the intended motion for animation [13]. Therefore, we introduce a new benchmark, TI2V-Bench, that consists of $320$ image-text pairs covering $22$ different scenarios. Each scenario includes $3$ to $5$ images with similar content in different styles, and $3$ to $5$ prompts to animate these images to generate videos with different motion. We include challenging scenarios such as novel object introductions and fine-grained object references. We use the Emu model [11], publicly available through meta.ai, to generate the initial images.
@@
Previous studies have used various automatic metrics to assess different quality aspects, such as video quality, motion intensity, and text alignment. We observed that these metrics may not always align with human perception. Following Movie Gen [30], we mainly rely on human evaluation with A-B testing to evaluate the performance. Our human evaluation protocol is motivated by the JUICE protocol [15], in which annotators are asked to indicate an overall preference and justify their choices across several axes, including image alignment, text alignment, object motion, and overall quality. Through our comprehensive human evaluations, we show that MotiF improves TI2V generation capabilities, notably improving text alignment and motion quality. In summary, we make the following contributions.
@@
  • We present Motion Focal Loss (MotiF) that encourages TI2V generation to concentrate on regions with larger motion and is complementary to existing techniques.
@@
  • We present a new benchmark, TI2V-Bench, including synthetic images and texts that cover a wide range of complex scenarios, as well as a human evaluation protocol designed to assess the performance of TI2V generation.
@@ 2.1 Text-Image-to-Video Generation
With the rapid advancements in Text-to-Image (T2I) and Text-to-Video (T2V) generation, TI2V generation has also gained significant attention. Some training-free methods [49, 28] utilize pretrained T2I or T2V models to enable zero-shot TI2V generation. Our review is focused on training-based approaches,  [1, 54, 37]. The techniques in prior TI2V generation methods can be mainly grouped into two categories: 1) design novel architectures to integrate the image condition, 2) derive motion priors to improve motion learning. Our approach falls into the second category.
Image Condition Integration. Most recent works use either the U-Net architecture [18, 17] or the Diffusion Transformer (DiT) [29] for video generation, where the image condition is incorporated similarly. Most works concatenate the image with the noisy video [15, 3, 25, 10], which acts as a spatial-aligned content guidance. Additionally, DynamiCrafter [46], ConsistI2V [33] and LivePhoto [10] also use cross-attention layers on the image embeddings to further strengthen the image condition and improve consistency in video generation. On the contrary, Zhao et al. [53] apply higher noise levels at later time steps to weaken the image condition to alleviate the condition image leakage issue. Follow-Your-Click [26] introduce a masking strategy to degrade the image condition. In MotiF, instead of disrupting the image condition to shift the model’s focus on the text, our training objective explicitly encourages the model to focus on video regions with more motion, and thus be more faithful to the text description.
Motion Priors. Our world exhibits a rich diversity of unique motion patterns. Modeling motion from real-world videos for generation is challenging due to the inherent ambiguity that arises when a single description can correspond to multiple possible motions. Many methods explore motion priors to improve motion learning. A common approach is to calculate a motion score [10, 12, 26, 51, 31, 25, 16] from the video and embed it into the model, which also enables motion intensity control during inference. LivePhoto [10] introduce a framework to re-weight text embeddings to place greater emphasis on the motion-related words. Cinemo [25] use a novel strategy to learn motion residuals. MoCA[48] leverage optical flow as an additional conditioning. Similar to our work, Follow-Your-Click [26] also use optical flow to calculate a mask, but the mask is embedded into the model and it requires additional input from the user during inference time. In our case, the motion prior is leveraged at the training objective level, which is explicit, simple and does not require any additional input during inference.
@@ 2.2 Video Generation Evaluation
Benchmarks and Automatic Metrics. Make-a-Video reports FVD/IS on the UCF-101 dataset [41] and FID/CLIP similarities on the MSR-VTT dataset [47] under the zero-shot setting, which is followed by later works in both T2V [4, 54, 1, 14] and TI2V generation [46, 33, 25, 53]. Re-purposing existing video understanding datasets for TI2V evaluation is undesirable due to their unconstrained data collection procedures, which may not capture the nuances of TI2V tasks. As a result, recent studies start to collect new evaluation sets [33, 13, 51] and explore additional automatic metrics [24, 21, 52]. AIGCBench [13] is proposed for TI2V task. However, their text descriptions are used to generate the image but not to drive the motion generation.AnimateBench [51] shares some similarities with our benchmark as using T2I models to generate images and applying different text prompts to animate the same image. However, the number of unique images and texts in AnimateBench is small and the domains are also limited.
Human Evaluation. Following Movie Gen [30], we rely on human annotators to evaluate the performance of TI2V generation. Prior works use different human evaluation protocols including A-B testing [53], A-testing [10] and ranking multiple methods [26, 16, 51]. While A-testing and ranking are more efficient, we believe A-B testing is less ambiguous and may yield more reliable results. Unlike previous works that evaluate multiple metrics, we ask the annotators to select the best video and justify their choices based on various aspects, following the JUICE design [15]. This will generate a single metric for calibration while also enabling a detailed examination of results through the various quality justifications. To the best of our knowledge, we have conducted the most comprehensive human evaluation by comparing to nine open-sourced TI2V models and showing the advantage of MotiF.
@@ 3.2 Modeling
Figure 3: Example image-text pairs in TI2V-Bench. For each scenario (column), we first think of a scene that could be potentially animated to generate different types of motion. We include challenging scenarios when there are multiple objects (yellow/blue/red balloon) in the initial image for fine-grained control or the text prompt describes a new object (frisbee, bubbles) to enter the scene. Then we come up with different prompts and use the publicly available meta.ai tool to generate diverse sets of images. Images of low quality or those not in the appropriate initial state are removed.
@@
TI2V-Bench (Ours) image(synthetic)-text Yes $320$ $88$ $133$ human TI2V score
@@ 4.1 Evaluation Set
While it’s relatively easy to collect a diverse set of prompts for T2V evaluation, evaluating TI2V is not straightforward as it requires a conditional image. As shown in Table 1, recent TI2V benchmarks can be classified into three categories: 1) video-text pairs; 2) image-text pairs with realistic images; 3) image-text pairs with synthetic images. Re-purposing existing video understanding datasets [41, 47, 2] for TI2V evaluation is non ideal, as these datasets assume only one possible outcome from the initial frame, and the text descriptions typically do not emphasize the intended motion. Existing images are also limited in their diversity and potential motion changes. Thus, we resort to image-text pairs with synthetic images to curate TI2V-Bench. AIGCBench [13] and Animate Bench [51] are most relevant to our work. However, the texts in AIGCBench are used to generate the images but not meant to describe the motion in the videos, which is critical for TI2V task. Animate Bench employs motion-based text descriptions, but its evaluation set is relatively small and limited, where the images often feature a single object or scene, and the text description lacks diversity.
Our goal is to curate an evaluation set that has a diverse set of images, and each image has a diverse set of prompts for text-guided animation. Some examples are shown in Figure 3. To create this benchmark, we first design scenarios in which multiple actions can be performed from an initial state and enumerate a range of possible actions. For example, in the scenario of “an aerial view of a car on the road," possible actions could include “moving forward," “moving backward," and “changing its lane." We come up with $22$ different scenarios with diverse objects and scenes, each with multiple options for different motion. We make sure to include challenging scenarios like images with multiple objects, and text descriptions that introduce a new object expected to appear in the video.
@@
Figure 4: Human evaluation results comparing MotiF to nine open-sourced models [46, 7, 25, 50, 9, 53, 28, 12, 33] on TI2V-Bench. We achieved considerable improvements across the board with an average preference of $72\%$. Through examining the justification choices, we found that our model mostly excel at improving text alignment and object motion, which matches very well with our motivation.
@@ 5 Experiment
Figure 5: Qualitative comparison to prior works on TI2V-Bench. Sampled frames are ordered from left to right.
@@ 5.2 Comparison to Prior Works
We compare to nine open-sourced TI2V generation methods.  [^1] Note that all the inference of prior works are done at Brown University. We ensure fair comparisons by following each method’s pre-processing and post-processing pipeline to ensure the input is in the optimal aspect ratio and the generated videos are without any undesirable distortion. The results are shown in Figure 4. Our method wins by a considerable margin to prior works. From the justification selections, the main reasons for our method to win are on the object motion and text alignment, which is exactly the motivation of MotiF to improve text-driven motion learning.
@@
In addition to human evaluation, we also report the automatic evaluation on Animate Bench [51] that is most relevant to TI2V-Bench. As shown in Table 4, MotiF achieved comparable performance to prior works. It’s worth noting that a static video baseline (by repeating the condition image) gets the best Image Alignment score and reasonable Text Alignment score, suggesting that these metrics are non ideal for evaluating TI2V generation. This has motivated us to mainly rely on human evaluation.